Sketch the graph of the parabola (y - 2)^2 = 12(x - 1) showing clearly the focus and the directrix
1
Expert's answer
2021-01-31T16:01:44-0500

"(y-k)^2=4p(x-h)"

focus: "(h+p,k)"

directrix: "x=h-p"


We have:

"k=2,p=3, h=1"

focus: "(4,2)"

directrix: "x=1-3=-2"